radical 2y-3 - radical y-3 = 1
Answer by nyc_function(2741)

sqrt{2y - 3} - sqrt{y - 3} = 1
Square both sides.
[sqrt{2y - 3} - sqrt{y - 3}]^2 = (1)^2
Can you take it from here?